OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations?

OpenStudy (anonymous):

3x + y = 14 –3x + 2y = 1

OpenStudy (aravindg):

add the 2 equations what do you get ?

OpenStudy (anonymous):

I am learning how to solve systems of equations by elimination.

OpenStudy (anonymous):

y=5

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

Will you explain? @Romero00

OpenStudy (anonymous):



OpenStudy (anonymous):

you have to add the equation together 3x+-3x is 0 so you add your ys y+2y =3y then you add your endings together 1+14=15

OpenStudy (anonymous):

then to get y you divide 3y/3 and 15/3 and your y=5

OpenStudy (anonymous):



OpenStudy (anonymous):

got it?